OpenStudy (anonymous):

I have no idea how to do division. Please help me. 5x^3y^4z / 10xy^2z^3

OpenStudy (anonymous):

there is some ambiguity in the denominator , please retype it , using proper bracketts

OpenStudy (jlastino):

When you divide exponents you simply subtract them x^3/x would be x^3-1 x^2 (but you can only divide x by x and y by y) I think you can do this problem having that in mind :D

OpenStudy (anonymous):

But when you do that what goes on top and what stays on the bottom?

OpenStudy (anonymous):

so it would be 2x^2y^2 over z^2?

OpenStudy (amistre64):

top - bottom; then flip the negatives afterwards

OpenStudy (anonymous):

okay thankies!!

OpenStudy (jlastino):

wait it should be 3y

OpenStudy (amistre64):

5/10 = 1/2 not 2 tho

OpenStudy (amistre64):

y^(4-2) = y^2

OpenStudy (anonymous):

wait im confused now...

OpenStudy (anonymous):

so then what would the answer be?

OpenStudy (amistre64):

you had it right; except for your constants; put the 2 under

OpenStudy (amistre64):

2x^2y^2 over z^2 ; bad x^2y^2 over 2z^2 ; gooder

OpenStudy (amistre64):

it would help out if you spaced between factors so that they dont all bleed together

OpenStudy (anonymous):

is your fraction \[\frac{5 x^3 y^4 z}{10 x y^2 z^3} \]

OpenStudy (anonymous):

if this is the case, the answer is \[\frac{x^2 y^2}{2 z^2} \]